How many moles of gas would be present in 0.23 L at STP?


1
Expert's answer
2020-12-14T14:49:29-0500

From PV=nRT

We have n=PV/RT

At stp P= 1atm T=273k R=8.314

n=1×0.23/8.314×273

n=1×10-4moles